OpenStudy (anonymous):

calculate the area of a circle with each radius,to the nearist square unit. A)5m B)12cm

OpenStudy (anonymous):

\[A = \pi r^{2}\]

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

I gave you the formula to find the answer. Just plug in the radius value and solve.

OpenStudy (amistre64):

that is the answer ...

OpenStudy (anonymous):

i dont know how though

OpenStudy (amistre64):

dont know how to do what? find squares or multiply or what exactly?

OpenStudy (anonymous):

find squares

OpenStudy (amistre64):

when given a number such as: 4 we can find the square by multiplying it to itself 4^2 = 4*4 = 16

OpenStudy (amistre64):

this concept applies to all squares; so, 5^2 = 5*5 = ?

OpenStudy (anonymous):

can you just do it and give me the answer

OpenStudy (amistre64):

this is the answer ... this is the part about studying that you interact with and gain a solid understanding of basic concepts

OpenStudy (amistre64):

or use a calculator :)

OpenStudy (anonymous):

what is the answer for 5m

OpenStudy (amistre64):

it is: pi r^2 ; when r=5

OpenStudy (anonymous):

whats the formul

OpenStudy (amistre64):

the formula for the area of a circle is: pi r^2

OpenStudy (anonymous):

for grade 8

OpenStudy (amistre64):

thats the formula for any grade; except maybe they say pi = 3.14 in grade 8; i cant recall. Its been quite awhile